# Spin-dependent terms of the Breit-Pauli Hamiltonian evaluated with explicitly correlated Gaussian basis set for molecular computations

### v1

@@ V.2 Fine structure of the $\text{b}\ ^{3}\Pi_{\text{g}}$ and $\text{c}\ ^{3}\Sigma_{\text{g}}^{+}$ states
:$\dagger$:† : $N_{\mathrm{B}}=500$ and $N_{\mathrm{C}}=750$.
::*:∗ : $R=2.015$ bohr bond distance for a ${}^{3}\Sigma_{u}^{+}$.
:$\circ$:∘ : $R=2.08$ bohr bond distance for c ${}^{3}\Sigma_{g}^{+}$.
@@
$N_{\mathrm{b}}$ $N_{\mathrm{c}}$ $\gamma^{(3)}_{2}$ $\delta^{(3)}_{1}$ $\delta^{(3)}_{2}$ $\lambda^{(3)}$ $\zeta^{(3)}$ $\xi^{(3)}$
@@
:$\dagger$:† : $N_{\mathrm{B}}=500$ and $N_{\mathrm{C}}=750$.
@@ VI Summary, conclusion, and outlook
For small molecular species, accurate and complete fine-structure data are scarcely available, and this work fills this gap. First of all, we tested the methodology for an atomic system, for which (higher precision) fine-structure splitting data are already available in the literature. Then, we computed spin-dependent Breit-Pauli Hamiltonian matrix elements for the two-electron triplet H2 and triplet H+${}_{3}^{+}$. High-precision literature data are not available for these fine-structure splittings; however, for these two-electron systems, we verified the results against our in-house developed no-pair Dirac-Coulomb-Breit (np-DCB) Hamiltonian approach, using the same fECG spatial basis set in the two computations. The two approaches give identical splittings to high precision (the most stringent, direct comparison would be possible by $\alpha$-scaling the np-DCB results and comparing the $\alpha^{2}E_{\text{h}}$ order term).
@@ Acknowledgement
PJ gratefully acknowledges the support of the János Bolyai Research Scholarship of the Hungarian Academy of Sciences (BO/285/22). We thank the European Research Council (Award No. 851421) and the Momentum Programme of the Hungarian Academy of Sciences (LP2024-15/2024) for their financial support. We acknowledge KIFÜ (Governmental Agency for IT Development, Hungary) for access to the Komondor HPC facility based in Hungary.
Supplemental Material
Spin-dependent terms of the Breit-Pauli Hamiltonian evaluated with explicitly correlated Gaussian basis set for molecular computations
@@
(Dated: 23 June 2025)
S1 Representation of the spin function with spin-adapted states and spinors - case study for four particles with $M_{S}=0$
@@
$|\Xi_{m_{s_{1}}m_{s_{2}}...m_{s_{n_{\text{el}}}}}\rangle=|\eta_{m_{s_{1}}}\rangle\otimes|\eta_{m_{s_{2}}}\rangle\otimes\dots\otimes|\eta_{m_{s_{n_{\mathrm{el}}}}}\rangle\,$ (S1)
where $|\eta_{m_{s_{i}}}\rangle$ is a single-particle spin function (vector). For $m_{s_{i}}=-1/2$, the spin function is ‘spin up’,
$|\eta_{0}\rangle=|\alpha\rangle=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\ ,$ (S4)
for $m_{s_{i}}=+1/2$, it is ‘spin down’,
$|\eta_{1}\rangle=|\beta\rangle=\left(\begin{array}[]{c}0\\ 1\end{array}\right)\ .$ (S7)
The $\{|\Xi_{m_{s_{1}}m_{s_{2}}...m_{s_{n_{\text{el}}}}}\rangle\}$ form a basis for spin eigenfunctions (see Ref. [65] and the Supplementary Material of Ref. 46). For four electrons with $M_{S}=0$, the spin-adapted states are as follows. We have two singlet ($S=0$) states,
$|\Theta_{0,0}^{(1)}\rangle$ $=\frac{1}{\sqrt{12}}\big{(}|\Xi_{0011}\rangle-2|\Xi_{0101}\rangle+|\Xi_{0110}\rangle+|\Xi_{1001}\rangle-2|\Xi_{1010}\rangle+|\Xi_{1100}\rangle\big{)}$
$|\Theta_{0,0}^{(2)}\rangle$ $=\frac{1}{2}\big{(}|\Xi_{0011}\rangle-|\Xi_{0110}\rangle-|\Xi_{1001}\rangle+|\Xi_{1100}\rangle\big{)}\ ,$
@@
$|\Theta_{1,0}^{(1)}\rangle$ $=\frac{1}{\sqrt{2}}\big{(}|\Xi_{0011}\rangle-|\Xi_{1100}\rangle\big{)}$
$|\Theta_{1,0}^{(2)}\rangle$ $=\frac{1}{\sqrt{2}}\big{(}|\Xi_{0101}\rangle-|\Xi_{1010}\rangle\big{)}$
$|\Theta_{1,0}^{(3)}\rangle$ $=\frac{1}{\sqrt{2}}\big{(}|\Xi_{0110}\rangle-|\Xi_{1001}\rangle\big{)}\ .$
and one set of quintuplet ($S=2$) state,
$|\Theta_{2,0}\rangle$ $=\frac{1}{\sqrt{6}}\left(|\Xi_{0011}\rangle+|\Xi_{0101}\rangle+|\Xi_{0110}\rangle+|\Xi_{1001}\rangle+|\Xi_{1010}\rangle+|\Xi_{1100}\rangle\right)\ .$
Despite this, the spin-adapted implementation becomes complicated for the general $N$-electron case, especially for the computation of matrix elements with spin-dependent relativistic operators. Therefore, in this work, we have switched to the spinor representation, in which the calculation of spin-dependent operator matrix elements can be more efficiently automated.
For the transformation between the two representations, i.e., the spin-adapted ‘$\boldsymbol{\theta}$’ parameterisation vs. spinor linear combination, the relation between $C_{k,S}$ and $d^{(SM_{S})}_{m_{s_{1}},...,m_{s_{n_{\text{el}}}}}$ is
$|\chi_{S,M_{S}}(\boldsymbol{\theta})\rangle$ $=\sum_{k=1}^{N_{\text{s}}(n_{\text{el}},S)}C_{k,S}(\boldsymbol{\theta})|\Theta^{(k)}_{S,M_{S}}\rangle\ ,$
$=\sum_{k=1}^{N_{\text{s}}(n_{\text{el}},S)}C_{k,S}(\boldsymbol{\theta})\sum_{\begin{subarray}{c}s_{1},...,s_{n_{\text{el}}}\\ \Sigma s_{i}=M_{S}\end{subarray}}Q^{S,M_{S},k}_{s_{1}...s_{n_{\text{el}}}}|\Xi_{s_{1}s_{2}...s_{n_{\text{el}}}}\rangle$
$=\sum_{\begin{subarray}{c}m_{s_{1}},...,m_{s_{n_{\text{el}}}}\\ \Sigma m_{s_{i}}=M_{S}\end{subarray}}d^{(SM_{S})}_{m_{s_{1}},...,m_{s_{n_{\text{el}}}}}(\boldsymbol{\theta})\,|\Xi_{s_{1}s_{2}...s_{n_{\text{el}}}}\rangle\ ,$ (S8)
and this change of representation is performed in an automated procedure in QUANTEN. For the example of the helium dimer’s case, the non-relativistic energy optimisation was performed in the spin-adapted $\boldsymbol{\theta}$ parameterization, which was converted to the spinor representation, which means storing a $2^{4}$-dimensional vector for every basis function (instead of the 1-2 $\vartheta$ parameters). The latter representation can be more straightforwardly used for the evaluation of the spin-dependent BP matrix elements.
Although the non-relativistic wave functions are degenerate in $M_{S}$, but we need to generate functions with another $M_{S}$ value for computing the BP matrix elements. For this purpose, Eq. (S8) and the matrix representation of the spin-ladder operators were employed.
@@
In the case of closely degenerate states such as the $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ and $\text{A}\ ^{1}\Sigma_{\text{u}}^{+}$ states, it is essential to treat them within a quasi-degenerate framework. This approach takes into account the small mixing of the two states due to the perturbation. For this purpose, we build up the cumulative second-order Hamiltonian, $\mathbf{H}^{[2]}$, which includes contributions from both the direct couplings and the second-order perturbative corrections,
@@
where the states $X$ and $Y$ are the states considered within the model space. The operator $\hat{Q}$ is the projector to orthogonal complement to the coupled ($X,Y$) subspace, e.g., $\hat{Q}=1-|X\rangle\langle X|-|Y\rangle\langle Y|$).
@@
Based on this formalism, we computed a numerical estimate for the second-order coupling of $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ and A ${}^{1}\Sigma_{\text{u}}^{+}$ through the f ${}^{3}\Pi_{\text{u}}$ and F ${}^{1}\Pi_{\text{u}}$ states, identified as the most important ‘interaction transmission’ partners based on Ref. [69]. Pilot computations suggest that the resulting second-order contribution is on the $10^{-12}\ E_{\text{h}}$ range, so completely negligible in this work.

### v2

@@ V.2 Fine structure of the $\text{b}\ ^{3}\Pi_{\text{g}}$ and $\text{c}\ ^{3}\Sigma_{\text{g}}^{+}$ states
$\dagger:$ $N_{\mathrm{B}}=500$ and $N_{\mathrm{C}}=750$.
:*: $R=2.015$ bohr bond distance for a $^{3}\Sigma_{\text{u}}^{+}$.
$\circ:$ $R=2.08$ bohr bond distance for c $^{3}\Sigma_{\text{g}}^{+}$.
@@
$N_{\mathrm{b}}$ $N_{\mathrm{c}}$ $\alpha^{3}\gamma^{(3)}_{2}$ $\alpha^{3}\delta^{(3)}_{1}$ $\alpha^{3}\delta^{(3)}_{2}$ $\alpha^{3}\lambda^{(3)}$ $\alpha^{3}\zeta^{(3)}$ $\alpha^{3}\xi^{(3)}$
@@
$\dagger:$ $N_{\mathrm{B}}=500$ and $N_{\mathrm{C}}=750$.
@@ VI Summary, conclusion, and outlook
For small molecular species, accurate and complete fine-structure data are scarcely available, and this work fills this gap. First of all, we tested the methodology for an atomic system, for which (higher precision) fine-structure splitting data are already available in the literature. Then, we computed spin-dependent Breit-Pauli Hamiltonian matrix elements for the two-electron triplet H2 and triplet $_{3}^{+}. H$igh-precision literature data are not available for these fine-structure splittings; however, for these two-electron systems, we verified the results against our in-house developed no-pair Dirac-Coulomb-Breit (np-DCB) Hamiltonian approach, using the same fECG spatial basis set in the two computations. The two approaches give identical splittings to high precision (the most stringent, direct comparison would be possible by $\alpha$-scaling the np-DCB results and comparing the $\alpha^{2}E_{\text{h}}$ order term).
@@ Acknowledgement
PJ gratefully acknowledges the support of the János Bolyai Research Scholarship of the Hungarian Academy of Sciences (BO/285/22). We thank the European Research Council (Grant No. 851421) and the Momentum Programme of the Hungarian Academy of Sciences (LP2024-15/2024) for their financial support. We acknowledge DKP for granting us access to the Komondor HPC facility.
Supporting Information
Spin-dependent terms of the Breit-Pauli Hamiltonian evaluated with an explicitly correlated Gaussian basis set for molecular computations
@@
(Dated: 27 August 2025)
S1 Representation of the spin function with spin-adapted states and spinors - case study for four electrons with $M_{S}=0$
@@
$|\Xi_{t_{1}t_{2}...t_{n_{\text{el}}}}\rangle=|\eta_{t_{1}}\rangle\otimes|\eta_{t_{2}}\rangle\otimes\dots\otimes|\eta_{t_{n_{\text{el}}}}\rangle\,$, (S1)
where $t_{i}=2m_{s_{i}}+1$ and $|\eta_{t_{i}}\rangle$ is a single-particle spin function (vector). For $t_{i}=1$, the spin function is ‘spin up’,
$|\eta_{1}\rangle=|\alpha\rangle=\left(\begin{array}[]{c}1\\ 0\end{array}\right)\ ,$ (S4)
for $t_{i}=0$, it is ‘spin down’,
$|\eta_{0}\rangle=|\beta\rangle=\left(\begin{array}[]{c}0\\ 1\end{array}\right)\ .$ (S7)
The $\{|\Xi_{t_{1}t_{2}...t_{n_{\text{el}}}};t_{i}=0,1\rangle\}$ set forms a basis set for the spin eigenfunctions (see Ref. 76 and the Supplementary Material of Ref. 26). For four electrons with $M_{S}=0$, the spin-adapted states are as follows (see also Ref. 9). We have two singlet ($S=0$) states,
$|\Theta_{0,0}^{(1)}\rangle$ $=\frac{1}{\sqrt{12}}\big{(}|\Xi_{0011}\rangle-2|\Xi_{0101}\rangle+|\Xi_{0110}\rangle+|\Xi_{1001}\rangle-2|\Xi_{1010}\rangle+|\Xi_{1100}\rangle\big{)}$ (S8)
$|\Theta_{0,0}^{(2)}\rangle$ $=\frac{1}{2}\big{(}|\Xi_{0011}\rangle-|\Xi_{0110}\rangle-|\Xi_{1001}\rangle+|\Xi_{1100}\rangle\big{)}\ ,$ (S9)
@@
$|\Theta_{1,0}^{(1)}\rangle$ $=\frac{1}{\sqrt{2}}\big{(}|\Xi_{1100}\rangle-|\Xi_{0011}\rangle\big{)}$ (S10)
$|\Theta_{1,0}^{(2)}\rangle$ $=\frac{1}{\sqrt{2}}\big{(}|\Xi_{1010}\rangle-|\Xi_{0101}\rangle\big{)}$ (S11)
$|\Theta_{1,0}^{(3)}\rangle$ $=\frac{1}{\sqrt{2}}\big{(}|\Xi_{1001}\rangle-|\Xi_{0110}\rangle\big{)}\ .$ (S12)
and one quintuplet ($S=2$) state,
$|\Theta_{2,0}\rangle$ $=\frac{1}{\sqrt{6}}\left(|\Xi_{0011}\rangle+|\Xi_{0101}\rangle+|\Xi_{0110}\rangle+|\Xi_{1001}\rangle+|\Xi_{1010}\rangle+|\Xi_{1100}\rangle\right)\ .$ (S13)
Instead of the computer implementation of these functions with abstract $|\alpha\rangle$ and $|\beta\rangle$ labels, we found it more convenient to use the vector representation for a general $N$-electron implementation, and especially for the computation of matrix elements with spin-dependent relativistic operators. Therefore, in this work, we have switched to the $2^{N}$-dimensional spinor representation.
For the transformation between the two representations, i.e., the spin-adapted $\theta$-parameterisation vs. spinor linear combination, the relation between $C_{k,S}$ and $d^{(SM_{S})}_{t_{1},...,t_{{n_{\text{el}}}}}$ is
$|\chi_{S,M_{S}}(\boldsymbol{\theta})\rangle$ $=\sum_{k=1}^{N_{\text{s}}(n_{\text{el}},S)}C_{k,S}(\boldsymbol{\theta})|\Theta^{(k)}_{S,M_{S}}\rangle\$,
$=\sum_{k=1}^{N_{\text{s}}(n_{\text{el}},S)}C_{k,S}(\boldsymbol{\theta})\sum_{\begin{subarray}{c}t_{1},...,t_{n_{\text{el}}}\\ \Sigma t_{i}=2M_{S}+1\end{subarray}}Q^{S,M_{S},k}_{t_{1}...t_{n_{\text{el}}}}|\Xi_{t_{1}t_{2}...t_{n_{\text{el}}}}\rangle$
$=\sum_{\begin{subarray}{c}t_{1},...,t_{{n_{\text{el}}}}\\ \Sigma t_{i}=2M_{S}+1\end{subarray}}d^{(SM_{S})}_{t_{1},...,t_{n_{\text{el}}}}(\boldsymbol{\theta})\,|\Xi_{t_{1}t_{2}...t_{n_{\text{el}}}}\rangle\ ,$ (S14)
and this change of representation is performed in an automated procedure in QUANTEN. For the example of the helium dimer’s case, the non-relativistic energy optimisation was performed in the spin-adapted $\theta$-parameterization [9, 64], which was converted to the spinor representation, which means storing a $2^{4}$-dimensional vector for every basis function (instead of 1-2 $\vartheta$ parameters of the $\theta$-parameterization, depending on the actual spin subspace). The latter representation can be more straightforwardly used to evaluate the spin-dependent Breit-Pauli (BP) matrix elements.
Although the non-relativistic wave functions are degenerate in $M_{S}$, we need to generate functions with another $M_{S}$ value for computing the BP matrix elements. For this purpose, Eq. (S14) and the matrix representation of the spin-ladder operators were employed.
@@
Closely degenerate states, such as the $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ and $\text{A}\ ^{1}\Sigma_{\text{u}}^{+}$ states of He2, can be treated in a quasi-degenerate framework. This approach takes into account the small mixing of the two states due to the perturbation. For this purpose, we build up the cumulative second-order Hamiltonian, $\mathbf{H}^{[2]}$, which includes contributions from both the direct couplings and the second-order perturbative corrections,
@@
where the states $X$ and $Y$ are the states considered within the model space. The operator $\hat{Q}$ is the projector to orthogonal complement to the coupled ($X,Y$) subspace, e.g., $\hat{Q}=1-|X\rangle\langle X|-|Y\rangle\langle Y|$.
@@
Based on this formalism, we computed a numerical estimate for the second-order coupling of $\text{a}\ ^{3}\Sigma_{\text{u}}^{+}$ and A $^{1}\Sigma_{\text{u}}^{+}$ through the f $^{3}\Pi_{\text{u}}$ and F $^{1}\Pi_{\text{u}}$ states, identified as the most important ‘interaction transmission’ partners based on Ref. [80]. Pilot computations suggest that the resulting second-order contribution is in the $10^{-12}\ E_{\text{h}}$ range, so completely negligible in this work.